OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point slope form, slope intercept form and standard form. Passing through (-2,-1) and perpendicular to the line whose equation is y+7=5/6(x-5)

OpenStudy (anonymous):

y+1=(-6/5)(x+2) y=(-6/5)x-17/5 6x+5y=-17

OpenStudy (anonymous):

perpendicular

OpenStudy (anonymous):

perpendicular slope is opposite and negative

OpenStudy (anonymous):

reciprocal and negative not opposite.

OpenStudy (anonymous):

I say opposite reciprocal since negative sometimes confuses younger students to think the perpendicular slope needs to be a negative number

OpenStudy (anonymous):

the 2 slopes multiply to give -1 is easiest, I think.